# Quantumness certification via non-demolition measurements

### v1

@@ VI Tracking the quantum-to-classical transition
A relevant application for the use of the QNDM approach is the tracking of the quantum-to-classical transition due to the interaction with the environment. The quantum-to-classical transition can be identified by looking at the reduction of the negative regions of $\mathcal{P}_{\rm QNDM}(\Delta)$ [22, 23].
As a prototypical example, let us consider a two-level system $S$ subject to an external drive, and interacting with an external engineered environment $E$. The engineered environment allows us to control the coupling strength between the system and the environment. The evolution of the system and the detector is as described above: a sequence of non-demolition measurements at discrete times $t_{i}$ via a fast, transient coupling with a detector $D$ interspersed with the system unitary transformations.
For the sake of clarity, here the physical problem is measuring the work, the internal energy variation, and the heat, as in Sec. III.1. But, since we are dealing with an open quantum system, the internal energy variation comprises both the work done on the system and the heat dissipated into the environment [22, 23]. To avoid any superfluous complexity, we will focus only on the internal energy variation, which is obtained by measuring the energy of the system at the start and the end of the thermodynamic transformation [22, 23].
@@
The Hamiltonian that governs the coupling between the system and the detector is $\hat{H}_{SD}(t)=f(t)\hat{H}_{S}(t)\otimes\hat{H}_{D}$, which is operated at times $t_{0}=0$ and $\mathcal{T}$. The system-detector coupling at these two times suffices to encode information about the internal energy variation of the system in the phase of the detector state. The time-dependence $f(t)$ in the coupling is chosen in order to implement the unitary transformations $\hat{u}_{0}=\exp\{-(i\lambda/\hbar)~\hat{\sigma}_{x}\otimes\hat{H}_{D}\}$ and $\hat{u}_{\mathcal{T}}=\exp\{-(i\lambda/\hbar)~\hat{\sigma}_{z}\otimes\hat{H}_{D}\}$. We assume that the system-detector coupling occurs over times much smaller than $\epsilon^{-1}$, which is the timescale of the system dynamics.
@@
where the first and the second qubits represent the system and the environment, respectively. The transformation in Eq. (VI) models the emission of an energy quantum with probability $p$, whereby we suppose that the environment can only induce relaxation, and not excitation in the system. The relaxation occurs in the basis that diagonalizes $\hat{H}_{S}(t)$, namely in the basis of $\hat{\sigma}_{x}$ for $0\leq t\leq t_{1}$ and in the basis of $\hat{\sigma}_{z}$ for $t_{1}\leq t\leq\mathcal{T}$. For simplicity, we denote with $\hat{R}_{x}$ and $\hat{R}_{z}$ the operators (acting on the system and environment degrees of freedom) that implement the transformation (VI). Since $\hat{R}_{x}$ and $\hat{R}_{z}$ generate two different dissipative dynamics, two distinct engineered environments must be implemented to simulate a Markovian environment: one engineered environment to realize $\hat{R}_{x}$ and the other for $\hat{R}_{z}$. Details on how this model can be implemented in a (commercial) superconducting quantum processor can be found in Ref. [22].
@@
The total system-detector-environment dynamics is described by the operator $\hat{U}_{\rm diss}=\hat{u}_{\mathcal{T}}~\hat{R}_{z}~\hat{U}_{z}~\hat{R}_{x}~\hat{U}_{x}~\hat{u}_{1}$. In other terms, first, we couple the system and the detector to measure the system energy at the initial time. Then, we let the system evolve with $\hat{U}_{x}$, and we implement the dissipative dynamics over the $\hat{\sigma}_{x}$ basis. Afterwards, the last sequence of operations is applied by changing the basis from $x$ to $z$. As we are considering a QNDM scheme, after the whole evolution, we measure the phase accumulated by the detector, and we repeat the procedure by using different couplings $\lambda$ in order to obtain the quasi-characteristic function $\mathcal{G}_{\lambda}$. Finally, the inverse Fourier transform is calculated to achieve the quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$.
@@ VII Conclusions
In this paper, we have reviewed the recent progress in interpreting and utilizing quantum non-demolition measurements to certify a violation of the macrorealism, specifically in relation to the presence of quantum interferences and coherent superpositions in the state of the quantum system under scrutiny.
The QNDM approach has two main advantages with respect to other protocols and tools present in the literature, such as LGIs [5, 27, 28, 30]. First, they give a necessary and sufficient condition for the macrorealism violation, which is directly linked with an operational procedure, discussed in Sec. IV.3. By applying such a procedure, indeed, one is certain to identify — if present — the presence of quantum coherence in the quantum system state. Second, the implementation of QNDM has been already successfully carried out in several experimental laboratories worldwide, as one can evince from Refs. [81, 82, 22, 83, 23, 84]. For these reasons, we believe that the approach based on QNDM is a privileged tool to identify quantum features of a system.
@@ VII.1 What’s Next?
  • Identification of a quantitative relation between the negativity of the QNDM quasi-probability distribution and a measure of quantum coherence for the current state of the measured system. The starting point would be Refs. [22, 23, 26]. Then, similar ideas might be extended to a measure of entanglement, in relation to local tests of Bell’s inequalities [18].
  • Detection of the quantum-to-classical transition, using the QNDM approach, for a quantum system approaching mesoscopic and even macroscopic scales. One of the open questions in the foundations of quantum mechanics, indeed, is to understand when a massive quantum body loses its quantum features, thus starting to behave classically. This line of research is still under investigation, as proved by Refs. [98, 99, 100, 101]. In particular, in the years, double-slit experiments have been performed using different bodies with an increasing mass. Since the QNDM tracks quantum interferences and coherent superpositions, it is worth testing experimentally possible violations of macrorealism, according to QNDM quasiprobabilities, in levitated mechanical systems of mesoscopic size, as in Ref. [101].
@@
  • Computation of quantum speed limit bound for non-demolition quasiprobabilities, taking inspiration from Ref. [108] that addresses this issue for Kirkwood-Dirac quasiprobabilities.
  • Extension of the proof of the sufficient and necessary condition for the violation of macrorealism to open quantum systems. In this regard, by setting ourselves in a context of monitoring heat generation, one could certify the presence of quantum contributions in the heat distribution using QNDM quasi-probabilities.

### v2

@@ VI Tracking the quantum-to-classical transition
A relevant application of the QNDM approach is the tracking of the quantum-to-classical transition due to the interaction with the environment. The quantum-to-classical transition can be identified by looking at the reduction of the negative regions of $\mathcal{P}_{\rm QNDM}(\Delta)$ [22, 23].
As a prototypical example, let us consider a two-level system $S$ subject to an external drive, and interacting with an external engineered environment $E$. The engineered environment allows us to control the coupling strength between the system and the environment. The evolution of the system and the detector is as described above: a sequence of non-demolition measurements at discrete times $t_{i}$ via a fast, transient coupling with a detector $D$ interspersed with unitary transformations applied to the system’s state.
For the sake of clarity, here the physical problem is measuring the work, the internal energy variation, and the heat. Since we are dealing with an open quantum system, the internal energy variation comprises both the work done on the system and the heat dissipated into the environment [22, 23]. To avoid any superfluous complexity, we focus only on the internal energy variation, which is obtained by measuring the system’s energy at the start and the end of the thermodynamic transformation [22, 23].
@@
The Hamiltonian that governs the coupling between the system and the detector is $\hat{H}_{SD}(t)=f(t)\hat{H}_{S}(t)\otimes\hat{H}_{D}$, which is operated at times $t_{0}=0$ and $\mathcal{T}$. The system-detector coupling at these two times suffices to encode information about the internal energy variation of the system in the phase of the detector state. The time-dependence $f(t)$ in the coupling is chosen in order to implement the unitary transformations $\hat{u}_{0}=\exp\{-(i\lambda/\hbar)~\hat{\sigma}_{x}\otimes\hat{H}_{D}\}$ and $\hat{u}_{\mathcal{T}}=\exp\{-(i\lambda/\hbar)~\hat{\sigma}_{z}\otimes\hat{H}_{D}\}$. We assume that the system-detector coupling occurs over times much smaller than $\epsilon^{-1}$, which is the timescale of the system’s dynamics.
@@
where the first and the second qubits represent the system and the environment, respectively. The transformation in Eq. (VI) models the emission of an energy quantum with probability $p$, whereby we suppose that the environment can only induce relaxation, and not excitation in the system. The relaxation occurs in the basis that diagonalizes $\hat{H}_{S}(t)$, namely in the basis of $\hat{\sigma}_{x}$ for $0\leq t\leq t_{1}$ and in the basis of $\hat{\sigma}_{z}$ for $t_{1}\leq t\leq\mathcal{T}$. For simplicity, we denote with $\hat{R}_{x}$ and $\hat{R}_{z}$ the operators that implement the transformation (VI). Since $\hat{R}_{x}$ and $\hat{R}_{z}$ generate two different dissipative dynamics, two distinct engineered environments must be implemented to simulate a Markovian environment: one engineered environment to realize $\hat{R}_{x}$ and the other for $\hat{R}_{z}$. Details on how this model can be implemented in a (commercial) superconducting quantum processor can be found in Ref. [22].
@@
The total system-detector-environment dynamics is described by the operator $\hat{U}_{\rm diss}=\hat{u}_{\mathcal{T}}\hat{R}_{z}\hat{U}_{z}\hat{R}_{x}\hat{U}_{x}\hat{u}_{1}$. In other terms, first, we couple the system and the detector to measure the energy of the system at the initial time. Then, we let the system evolve with $\hat{U}_{x}$, and we implement the dissipative dynamics over the $\hat{\sigma}_{x}$ basis. Afterwards, the last sequence of operations is applied by changing the basis from $x$ to $z$. As we are considering a QNDM scheme, after the whole evolution, we measure the phase accumulated by the detector, and we repeat the procedure by using different couplings $\lambda$ in order to obtain the quasi-characteristic function $\mathcal{G}_{\lambda}$. Finally, the inverse Fourier transform is calculated to achieve the quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$.
@@ VII Conclusions
In this paper, we have reviewed the recent progress in interpreting and utilizing quantum non-demolition measurements to certify the presence of quantum interferences and coherent superpositions in the state of the quantum system under scrutiny.
The QNDM approach has two main advantages with respect to other protocols and tools present in the literature, such as LGIs [5, 86, 87, 84]. First, they give a necessary and sufficient condition for the violation of the macrorealism per se, which is directly linked with an operational procedure, discussed in Sec. IV.2. By applying such a procedure, indeed, one is certain to identify — if present — quantum coherence in the quantum system’s state, independently of the initial state, measurement observables, or time evolution of the system. These violations are certified by examining a single quantity: the negativity of the QNDM quasi-probability density function. Second, QNDM have been successfully implemented in many experimental laboratories worldwide, as one can evince from Refs. [93, 94, 22, 95, 23, 96]. This is because the QNDM approach only requires coupling the measured system with a detector via a phase operator that has broad experimental applicability. In particular, in Refs. [22, 23], QNDM have been employed to monitor the loss of coherence due to environmental effects in commercial IBM quantum processors. These tests have shown that a QNDM protocol can be realized even in a setting hardly affected by noise. For these reasons, we believe that the approach based on QNDM is a privileged tool to identify quantum features of a system.
@@ VII.1 What’s Next?
  • Identification of a quantitative relation between the negativity of the QNDM quasi-probability density function and a measure of quantum coherence for the current state of the measured system. The starting point would be Refs. [22, 23, 26]. Then, similar ideas might be extended to a measure of entanglement, in relation to local tests of Bell’s inequalities [18].
  • Detection of the quantum-to-classical transition, using the QNDM approach, for a quantum system approaching mesoscopic and even macroscopic scales. One of the open questions in the foundations of quantum mechanics, indeed, is to understand when a massive quantum body loses its quantum features, thus starting to behave classically. This line of research is still under investigation, as proved by Refs. [97, 98, 99, 100, 101]. In particular, in recent years, double-slit experiments have been performed using different bodies with an increasing mass. Since the QNDM tracks quantum interferences and coherent superpositions, it is worth testing experimentally possible violations of macrorealism per se, according to QNDM quasi-probabilities, in levitated mechanical systems of mesoscopic size, as in Ref. [100].
  • The QNDM protocol resembles the decoherent histories approach to quantum theory, introduced by Gell-Mann, Hartle, Griffiths, Omnes in the mid-nineties [102, 103, 104, 105, 106, 107, 108] and discussed recently in Refs. [109, 110, 111]. Indeed, the decoherence functional in these cited papers comprises exactly the interference contributions in Eq. (14) for the quantum system to evolve along superposed trajectories. Moreover, the classical behaviour of the system under scrutiny obtained when the QNDM quasi-probability density function is positive [see Eq. (24)] corresponds to the decoherence functional going to zero. In this perspective, the QNDM protocol could be used to experimentally test the results and predictions of the decoherent histories approach. However, we must stress that the number of possible quantum trajectories to be tracked during the dynamics increases rapidly with the number of measurements performed. This limits the applicability of QNDM, because the quasi-probability spectrum soon becomes too complex to disentangle the contributions of the different trajectories.
@@
  • Computation of quantum speed limit bound for non-demolition quasi-probabilities, taking inspiration from Ref. [118] that addresses this issue for Kirkwood-Dirac quasi-probabilities.
  • Extension of the proof of the sufficient and necessary condition for the violation of the macrorealism per se to open quantum systems. In this regard, by setting ourselves in a context of monitoring heat generation, one could certify the presence of quantum contributions in the heat distribution using QNDM quasi-probabilities.